Which is expected to have the largest dispersion forces? Which is expected to have the largest dispersion forces? C18H36 K NH3 C4H8
Answer #1

Dispersion forces depend on the magnitude of molar mass of molecule.

The one with greater molar mass will have greater dispersion forces.

Molar mass of C18H36,

MM = 18*MM(C) + 36*MM(H)

= 18*12.01 + 36*1.008

= 252.468 g/mol

Molar mass of K = 39.1 g/mol

Molar mass of NH3,

MM = 1*MM(N) + 3*MM(H)

= 1*14.01 + 3*1.008

= 17.034 g/mol

Molar mass of C4H8,

MM = 4*MM(C) + 8*MM(H)

= 4*12.01 + 8*1.008

= 56.104 g/mol

Molar mass is largest for C18H36.

So, C18H36 has largest dispersion forces.

Answer: C18H36